OpenStudy (anonymous):

solve please 2/3x-2=5/6

OpenStudy (helder_edwin):

\[ \large \frac{2}{3}x-2=\frac{5}{6} \] right??

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

\[ \large \frac{2}{3}x=\frac{5}{6}+2=\frac{17}{6} \] \[ \large x=\frac{17}{6}\cdot\frac{3}{2}=\frac{17}{4} \]